40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Mar 10 16:29:32 2016, gautam, Update, endtable upgrade, Inventory check Inventory_check.pdfDoubler_comparison.pdf
    Reply  Thu Mar 10 16:32:45 2016, rana, Update, endtable upgrade, Inventory check 
       Reply  Mon Mar 14 22:42:23 2016, gautam, Update, endtable upgrade, Inventory check 
       Reply  Mon Mar 28 10:43:18 2016, Steve, Update, endtable upgrade, ETMX 4'x2' optical table pictures ETMX4x2Layout.jpgETMX_4x2.JPGETMXopt_4x2.JPG
          Reply  Wed Mar 30 17:38:52 2016, gautam, Update, endtable upgrade, X end table proposed layout  ETMX_3x2.JPGlayout_details_20160328.pdfETMX_proposed_layout.pdf
             Reply  Thu Mar 31 09:38:41 2016, Steve, Update, endtable upgrade, Y end 4x3 existing layout ETMY4x3layout.jpg
                Reply  Thu Mar 31 19:49:31 2016, gautam, Update, endtable upgrade, proposed layout v2 ETMX_proposed_layout_v2.pdf
                   Reply  Mon Apr 4 10:59:12 2016, gautam, Update, endtable upgrade, proposed layout v3 ETMX_proposed_layout_v3.pdfIR_modematch.pdfGreen_modematch.pdfXendALaMode.zipETMX_proposed_layout_dimensions.dwg
                      Reply  Mon Apr 4 15:04:14 2016, gautam, Update, endtable upgrade, COMPONENT REMOVAL 
                         Reply  Tue Apr 5 11:42:17 2016, gaericqutam, Update, endtable upgrade, TABLE REMOVAL 
                            Reply  Wed Apr 6 17:52:21 2016, gautam, Update, endtable upgrade, First contact cleaning commenced 
                               Reply  Thu Apr 7 12:51:24 2016, gautam, Update, endtable upgrade, Beam height differences 
                                  Reply  Thu Apr 7 14:30:58 2016, Steve, Update, endtable upgrade, ETMX table height ETMXbrs.pngETMX.jpgIMG_0618.JPGETMX.jpgETMXwc.jpg
                                     Reply  Fri Apr 8 09:13:20 2016, Steve, Update, endtable upgrade, ETMX-T beam height ETMXalignedT.pngETMXF.jpg
Message ID: 12065     Entry time: Wed Apr 6 17:52:21 2016     In reply to: 12063     Reply to this: 12066
Author: gautam 
Type: Update 
Category: endtable upgrade 
Subject: First contact cleaning commenced 

I've begun cleaning the optics that will eventually go back onto the newly installed X-endtable. We decided that First Contact was the way to go (as opposed to methanol drag wiping). Koji demonstrated the application of the (red) First Contact solution onto a 2" mirror - I then proceeded to work on the rest of the optics. We are broadly following the procedure in E1000079 - first one coat of First Contact solution is applied, then a small piece of PEEK is embedded by applying a second layer of solution over it (this will enable us to pull off the First Contact once we are ready - the plan is to do this after roughly placing the optic on the table. As of now, I've finished coating most of the optics that are part of the IR Transmon path - I will continue later in the evening.

The new endtable is almost ready for re-population. Steve just needs to shim the enclosure which will be done tomorrow morning. The game-plan as discussed at the meeting today is to first try and set up the IR Transmon path. This will allow us to verify that the endtable height is such that we can maintain a beam height of 4" everywhere on the table (I suspect we may have to compromise at some poing and do some fine adjustment of 1/4 to 1/2" somewhere though). It will also allow me to define the cavity axis relative to the table, which will be useful to place the green steering optics eventually. Doing this will be challenging though as right now, I can't see any of the arm flashes on the endtable using an IR card. Ideally, we want to somehow lock the X arm and then do the checks mentioned at the endtable, before beginning to put the endtable back together. 

ELOG V3.1.3-